If the volume of a gas container at 20.3oC changes from 3.1L to 9,172mL, what will the final temperature be (use units of Kelvin)? Assume the pressure doesn’t change and the amount of gas in the container doesn’t change. Remember that there are 1,000 mL in 1 liter.
Answer #1

At constant pressure and constant number of moles volume of a gas is directly proportional to Temperature of tha gas. and it is charle's law.

V1/T1 = V2/T2

3.1 * 1000 / (20.3+273) = 9172/T2

T2 = 867.789 K

T2 = 867.789 - 273 = 84.789 degree C